# Privacy-Preserving Resilient Vector Consensus

### v1

@@ 7 Conclusion
In this paper, we investigated private resilient vector consensus by adding Gaussian noise for a multi-agent system where there might be faulty agents. We first showed the resilient vector consensus for the normal agents in expectation. For the convergence accuracy, We proposed to analyze from the perspective of Mahalanobis distance between the final value and its expectation and Hausdorff distance between the two convex hulls with and without noise, respectively. Then we proved the $\rho-$concentrated geo-privacy for their initial states, and compare it with $(\varepsilon,\delta)-$differential privacy. At last, we demonstrated our result through numerical simulations. In the future, we will figure out the detailed distribution of the final value instead of using Chebyshev’s inequality. Moreover, it is desirable to compare the performance of algorithms with Laplace noise and Gaussian noise thoroughly.

### v2

@@ 7 Conclusion
In this paper, we studied private resilient vector consensus in multi-agent systems with faulty agents by adding Gaussian noise. We established resilient vector consensus in expectation and analyzed convergence accuracy using the Mahalanobis distance between the final value and its expectation and the Hausdorff distance between the two convex hulls with and without noise. We then proved $\rho$-CGP for initial states and compared it with $(\varepsilon,\delta)$-differential privacy. Numerical simulations validated our results. Future work will focus on identifying necessary and sufficient conditions for resilient consensus, finding the explicit lower bound for the nonzero entr1ies of $\mathbf{M}(t)$, deriving the exact distribution of the final value, and optimizing noise selection for CGP under different scenarios.